# State of the Art in Text Classification for South Slavic Languages: Fine-Tuning or Prompting?

### v1

@@ A.2.2.  Instruction-Tuned Large Language Models
  • OpenAI models, namely the GPT-3.5-Turbo (gpt-3.5-turbo-0125) (openai), GPT-4o (gpt-4o-2024-08-06) (openai-gpt4o) and the GPT-5 (gpt-5-2025-08-07) (gpt-5). These closed-source instruction-tuned LLMs were developed by OpenAI. OpenAI states that the models are trained on large multilingual web corpora, however, specific details about the training data, procedures, and architecture are not publicly known.
@@
  • Gemini 2.5 Flash model (comanici2025gemini) is a closed-source multilingual and multimodal instruction-tuned LLM by Google DeepMind. The model is reported to be pretrained on over 400 languages (comanici2025gemini), however, details on the language coverage are not available.
@@
  • Mistral Medium 3.1 model (mistral-medium-2508) (mistral) is a closed-source multimodal instruction-tuned model by Mistral AI. Available details on the model architecture and language coverage are very limited.
@@
  • LLaMA 3.3 model [^14] https://ollama.com/library/llama3.3 (llama33modelcard) is an open-source instruction-tuned multilingual LLM, developed by Meta, with 70 billion parameters. The model was pretrained on a web text collection in various languages, however, it is reported to support only 8 languages, namely, English, German, French, Italian, Portuguese, Hindi, Spanish, and Thai.
@@
  • Gemma 3 model [^15] https://ollama.com/library/gemma3 (gemma3) is an open-source multilingual instruction-tuned LLM, developed by Google DeepMind. The model was pretrained on multimodal data with large quantities of multilingual texts and is reported to support over 140 languages. We use the model in 27 billion parameter size.
@@
  • DeepSeek-R1-Distill [^16] https://ollama.com/library/deepseek-r1:14b (guo2025deepseek) is an open-source reasoning LLM, developed by DeepSeek AI. We use the distilled model in 14 billion parameter size, namely the DeepSeek-R1-Distill-Qwen-14B model. The model is based on the Qwen 2.5 model (qwen2.5; team2024qwen2) that was fine-tuned using a dataset curated with the DeepSeek-R1 reasoning model. The Qwen 2.5 model provides multilingual support for over 29 languages, including Chinese, English, French, Spanish, Portuguese, German, Italian, Russian, Japanese, Korean, Vietnamese, Thai, and Arabic.
@@
  • Qwen 3171717https://ollama.com/library/qwen3 (Qwen3-2504) (yang2025qwen3) is an open-source LLM, developed by Alibaba Cloud. We use the model with the 32 billion parameter size, namely, the qwen3:32b model. The model is said to support over 100 languages and dialects (yang2025qwen3).
@@
Open-source models were installed locally and executed via the Ollama API service (maric_local_llms). We use the quantized versions of the models as they are available through the Ollama library. [^18] https://ollama.com/library OpenAI models are used through the chat completion endpoint via the OpenAI API, whereas other closed-source models were accessed through the OpenRouter platform [^19] https://openrouter.ai/ that provides a unified API access to various closed-source models.
@@
To prevent any bias, all models were used with their default parameters. The only parameter that we defined is the temperature which we set to 0 to ensure a more deterministic behaviour of the models. The same prompts were used for all open-source and closed-source models. In Figure 4, we provide prompts that were provided to the LLMs for zero-shot text classification, namely for sentiment classification (Figure 4(a)), automatic genre identification (Figure 4(b)), news topic classification (Figure 4(c)) and topic classification in parliamentary speeches (Figure 4(d)). For more details on the setups used to apply fine-tuned BERT-like models and instruction-tuned LLMs to the test datasets, refer to the code published on GitHub. [^20] https://github.com/TajaKuzman/Benchmarking-Text-Classification-on-South-Slavic

### v2

@@ A.2.2.  Instruction-Tuned Large Language Models
  • OpenAI models, namely the GPT-3.5-Turbo (gpt-3.5-turbo-0125) (OpenAI, 2023), GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2024) and the GPT-5 (gpt-5-2025-08-07) (OpenAI, 2025). These closed-source instruction-tuned LLMs were developed by OpenAI. OpenAI states that the models are trained on large multilingual web corpora, however, specific details about the training data, procedures, and architecture are not publicly known.
@@
  • Gemini 2.5 Flash model (Comanici et al., 2025) is a closed-source multilingual and multimodal instruction-tuned LLM by Google DeepMind. The model is reported to be pretrained on over 400 languages (Comanici et al., 2025), however, details on the language coverage are not available.
@@
  • Mistral Medium 3.1 model (mistral-medium-2508) (Mistral AI, 2025) is a closed-source multimodal instruction-tuned model by Mistral AI. Available details on the model architecture and language coverage are very limited.
@@
  • LLaMA 3.3 model [^16] https://ollama.com/library/llama3.3 (Meta, 2024) is an open-weight instruction-tuned multilingual LLM, developed by Meta, with 70 billion parameters. The model was pretrained on a web text collection in various languages, however, it is reported to support only 8 languages, namely, English, German, French, Italian, Portuguese, Hindi, Spanish, and Thai.
@@
  • Gemma 3 model [^17] https://ollama.com/library/gemma3 (Gemma Team et al., 2025) is an open-weight multilingual instruction-tuned LLM, developed by Google DeepMind. The model was pretrained on multimodal data with large quantities of multilingual texts and is reported to support over 140 languages. We use the model in 27 billion parameter size.
@@
  • DeepSeek-R1-Distill [^18] https://ollama.com/library/deepseek-r1:14b (Guo et al., 2025) is an open-weight reasoning LLM, developed by DeepSeek AI. We use the distilled model in 14 billion parameter size, namely the DeepSeek-R1-Distill-Qwen-14B model. The model is based on the Qwen 2.5 model (Qwen Team, 2024b, a) that was fine-tuned using a dataset curated with the DeepSeek-R1 reasoning model. The Qwen 2.5 model provides multilingual support for over 29 languages, including Chinese, English, French, Spanish, Portuguese, German, Italian, Russian, Japanese, Korean, Vietnamese, Thai, and Arabic.
@@
  • Qwen 3191919https://ollama.com/library/qwen3 (Qwen3-2504) (Yang et al., 2025) is an open-weight LLM, developed by Alibaba Cloud. We use the model with the 32 billion parameter size, namely, the qwen3:32b model. The model is said to support over 100 languages and dialects (Yang et al., 2025).
@@
Open-weight models were installed locally and executed via the Ollama API service (Marić et al., 2025). We use the quantized versions of the models as they are available through the Ollama library. [^20] https://ollama.com/library OpenAI models are used through the chat completion endpoint via the OpenAI API, whereas other closed-source models were accessed through the OpenRouter platform [^21] https://openrouter.ai/ that provides a unified API access to various closed-source models.
@@
To prevent any bias, all models were used with their default parameters. The only parameter that we defined is the temperature which we set to 0 to ensure a more deterministic behaviour of the models. The same prompts were used for all open-weight and closed-source models. In Figure 4, we provide prompts that were provided to the LLMs for zero-shot text classification, namely for sentiment classification (Figure 4(a)), automatic genre identification (Figure 4(b)), news topic classification (Figure 4(c)) and topic classification in parliamentary speeches (Figure 4(d)). For more details on the setups used to apply fine-tuned BERT-like models and instruction-tuned LLMs to the test datasets, refer to the code published on GitHub. [^22] https://github.com/TajaKuzman/Benchmarking-Text-Classification-on-South-Slavic